OpenStudy (anonymous):

a printer charges $2.55 for binding plus 5cents per page for each copy of a community groups annual report. what is the fixed cost? What is the variable cost? let y represent the cost of the annual report and x represent the number of pages. what is the cost equation? y =___

OpenStudy (anonymous):

help

OpenStudy (a_clan):

y=$2.55 + 5cents*x convert $ into cents

OpenStudy (anonymous):

fixed cost is the binding charge of $2.55

OpenStudy (anonymous):

variable cost is the per page cost as it varies according to the number of pages

OpenStudy (anonymous):

y = 2.55 + 0.05x 5 cents = 5/100 $ = 0.05$

OpenStudy (anonymous):

fixed cost is 2.55$ and let number of pages =x variable cost is 5/100* X=.05X